OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 4x2 + 6y2 = 24. @rajee_sam @jim_thompson5910

OpenStudy (rajee_sam):

if you rewrite it in std. form you will have x²/6 + y²/4 = 1

OpenStudy (rajee_sam):

now my center is (h,k) which is (0,0) a² = 6 , a = √6 b² = 4 , b = 2 for ellipse a² - c² = b² c² = a² - b² = 6-4 = 2 c = √2 Vertices : ( h± a , k) , ( 0 + √6 , 0) and (0-√6 , 0) (√6, 0) and ( -√6 , 0) Foci : ( h ± c , k), ( 0 + √2 , 0) and ( 0-√2 , 0) (√2,0) and (-√2 , 0)

OpenStudy (anonymous):

those aren't any of my answer choices? @rajee_sam

OpenStudy (rajee_sam):

what are your choices?

OpenStudy (anonymous):

oh wait nevermind my bad it is

OpenStudy (anonymous):

could you help with one more? you're a lifesaver dude thanks so much! @rajee_sam

OpenStudy (rajee_sam):

I am not dude

OpenStudy (rajee_sam):

But I 'll help you never the less

OpenStudy (rajee_sam):

Are you understanding how to do it?

OpenStudy (anonymous):

Eliminate the parameter. x =1/2t, y = 3t^3 - 1 Oh my gosh i'm so sorry!

OpenStudy (anonymous):

i think it's y = 24x^3 - 1 but i'm not 100% sure

OpenStudy (rajee_sam):

for eliminating parameters you have to eliminate the t and express y in terms of x. for that you have to find what t is in terms of x. So use x = 1/2t and solve for t. Can you tell me what t is from x = 1/2t?

OpenStudy (anonymous):

so i'm guessing y = 24x^3 - 1 is not the right answer?

OpenStudy (rajee_sam):

no

OpenStudy (anonymous):

i'm stuck. I have no idea what you mean

OpenStudy (rajee_sam):

you are given 2 equations x = 1/2t ----(Equation 1) and y = 3t³ -1 ----(equation 2)

OpenStudy (rajee_sam):

In equation 1 you have x on one side and t on the other side

OpenStudy (rajee_sam):

you have to solve for t meaning you have to rewrite this equation as t = "something" Can you do that?

OpenStudy (anonymous):

t = 2?

OpenStudy (rajee_sam):

\[x = \frac{ 1 }{ 2t }\] If I rearrange the x and t from this equation I get , t = 1/2x

OpenStudy (rajee_sam):

Now I know what t is in terms of x

OpenStudy (rajee_sam):

put this in place of t in equation no. 2

OpenStudy (rajee_sam):

\[y = 3 (\frac{ 1 }{ 2x })^{3} - 1\]

OpenStudy (rajee_sam):

\[y = 3(\frac{ 1 }{ 8x ^{3} }) - 1\]\[y = \frac{ 3 }{ 8x ^{3} } - 1\] Now I have eliminated the parameter

OpenStudy (rajee_sam):

@hfskibo

OpenStudy (anonymous):

ohh ok that makes so much sense. Thanks! can you hellp with another one?

OpenStudy (rajee_sam):

last one I have to go after that

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (rajee_sam):

I hope you are learning in the process and not just getting answers out of me

OpenStudy (rajee_sam):

I would suggest you try similar problems without help and familiarize yourselves with the conic sections.

OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates.

OpenStudy (anonymous):

I am learning! you're doing a great job of explaining & i understand all these so much better now! thank you!

OpenStudy (anonymous):

@rajee_sam

OpenStudy (rajee_sam):

Ok

OpenStudy (rajee_sam):

You are given a polar coordinate which is ( r, Θ) ---> ( 3, 2π/3) r = 3 and Θ = 2π/3 = 120° Now rectangular coordinates are written as (x,y) your usual point coordinates. now this is calculated as follows using your polar coordinates x = r cos Θ y = r sin Θ you know r and Θ, so find your x and y and write your coordinates.

OpenStudy (rajee_sam):



OpenStudy (rajee_sam):

Can you find the coordinates?

OpenStudy (anonymous):

-3,2?